# Undersmoothed LASSO Models for Propensity Score Weighting and Synthetic Negative Control Exposures for Bias Detection

### v2

@@ Abstract
The propensity score (PS) is often used to control for large numbers of covariates in high-dimensional healthcare database studies. The least absolute shrinkage and selection operator (LASSO) is a data-adaptive prediction algorithm that has become the most widely used tool for large-scale PS estimation in these settings. However, recent work has shown that the use of data-adaptive algorithms for PS estimation can come at the cost of slow convergence rates, resulting in PS-based causal estimators having poor statistical properties. While this can create challenges for the use of data-driven algorithms for PS analyses, both theory and simulations have shown that LASSO PS models can converge at a fast enough rate to provide asymptotically efficient PS weighted causal estimators. In order to achieve asymptotic efficiency, however, LASSO PS weighted estimators need to be properly tuned, which requires undersmoothing the fitted LASSO model. In this paper, we discuss challenges in determining how to undersmooth LASSO models for PS weighting and consider the use of balance diagnostics to select the degree of undersmoothing. Because no tuning criteria is universally best, we propose using synthetically generated negative control exposure studies to detect bias across alternative analytic choices. Specifically, we show that synthetic negative control exposures can identify undersmoothing techniques that likely violate partial exchangeability due to lack of control for measured confounding. We use a series of numerical studies to investigate the performance of alternative balance criteria to undersmooth LASSO PS-weighted estimators, and the use of synthetic negative control exposure studies to detect biased analyses.
@@ 1 Introduction
Routinely collected healthcare data, including administrative claims and electronic health records, are increasingly being used to generate real‐world evidence on the effects of medical products. However, confounding stemming from nonrandomized exposures remains a fundamental obstacle to effectively utilizing these data sources for real-world evidence generation. To improve confounding control in healthcare database studies, data‐driven algorithms can be used to leverage the large volume of information in these data sources to identify features that indirectly capture information on unspecified confounding factors [1, 2, 3, 4]. A large literature has shown that supplementing investigator-specified covariates with large numbers of empirically identified features can often improve confounding control compared with adjustment for investigator-specified covariates alone [1, 2, 3, 4, 5]. In healthcare database studies, however, outcome events are often rare, making it difficult to empirically identify and model outcome associations for large numbers of covariates. Consequently, the propensity score (PS), defined as the conditional probability of exposure given a set of covariates, is widely used for large-scale covariate adjustment in these settings [6, 7].
Fitting large-scale PS models can capture more confounding factors, but it can also result in reduced covariate overlap, which can harm the properties of effect estimates and even violate positivity assumptions [8, 9]. Consequently, when fitting large-scale PS models, covariate selection is necessary to avoid problems of nonoverlap. The least absolute shrinkage and selection operator (LASSO) is a data-adaptive prediction algorithm that has become the most widely used tool for fitting large-scale PS models [4, 10]. LASSO uses penalized regression to data-adaptively shrink imprecise coefficients toward zero, effectively performing covariate selection to exclude less important predictors. However, recent work has shown that the use of data-adaptive machine-learning algorithms for estimating the PS can come at the cost of slow convergence rates. Specifically, it has been shown that if the PS model does not converge at a minimum rate of $n^{-1/4}$ (where $n$ is the sample size), then this can result in PS-based causal estimators having poor statistical properties with performance deteriorating as the dimension of the data increases (the curse of dimensionality) [11, 12, 13].
In recent work, Benkeser & van der Laan [14] showed that if the true regression function can be expressed as a linear combination of some expanded basis of the covariates (discussed further in Section 2), then a LASSO regression that is fitted on the expanded set of basis functions can estimate the true model at a rate of $n^{-1/3}$. Benkesser & van der Laan [14] further explain that since a very general class of functions can be expressed as a linear combination of an expanded set of indicator basis functions, LASSO regression is generally applicable for estimating nuisance functions at a fast enough rate for causal estimators to have desirable properties. Building on this work, Ertefaie et al [15] showed that LASSO models can provide asymptotically linear PS weighted estimators with variance converging to the nonparametric efficiency bound.
@@
To achieve asymptotic efficiency, however, LASSO PS weighted estimators need to be properly tuned, which requires undersmoothing the fitted LASSO model (discussed further in Section 2). Ertefaie et al [15] showed that the degree of undersmoothing can be derived from the form of the target causal parameter’s efficient influence function, if that function is available. In practice, however, the efficient influence function is often not known or difficult to derive. Even when the analytic form of the efficient influence function is known, using the efficient influence function to undersmooth the LASSO model requires modeling the conditional outcome mean [15], which can be difficult in studies with rare outcome events. It remains unclear how to best tune LASSO PS weighted estimators when the efficient influence function is difficult to derive or unknown, or when modeling the conditional outcome mean is challenging.
@@
In this work, we discuss challenges in determining how to undersmooth LASSO models for PS weighting and consider the use of balance diagnostics to select the degree of undersmoothing. Because no tuning criteria is universally best, we propose a framework to generate synthetic negative control exposure studies for bias detection. We show that if a given LASSO PS weighted analysis does not result in conditional independence between the synthetic exposures and observed outcome within unexposed individuals, then the same analysis is unlikely to satisfy partial exchangeability when applied to the original study population. Finally, we use a series of numerical studies to evaluate the performance of alternative undersmoothing criteria and the use of synthetic negative control exposure studies to detect biased analyses.
@@ 2.1 Framework & Setup
We assume that exposure is conditionally exchangeable given $X$, written as
@@
[…]
@@
We further assume positivity and consistency [19]. Positivity, formally written as $10<e(X)<$ 1, implies that the true PS function, $e(X)$, is bounded away from 0 and 1. Consistency implies that the observed outcome for each individual is equal to the potential outcome under their observed exposure status, written as $Y^{a}=Y$ for $A=a$.
In addition to the standard assumptions for causal inference described above, we assume that the logit of the propensity score function, $logit(e(X))$, is linear in $X$, written as:
$logit(e(X))=\beta_{0}+X^{\top}\beta$ (1)
where $\beta$ is a $d$-dimensional vector of parameters and $\beta_{0}$ is a scalar. Healthcare database studies often consist of adjusting for a high-dimensional set sparse binary covariates where linearity assumptions in the PS model are often reasonable [1, 2, 3, 4, 5]. When the assumption of linearity in the relationship between $X$ and the logit of the propensity score is not reasonable, Benkeser & van der Laan [14] showed that under mild global smoothness assumptions, the baseline covariates, $X$, can be expanded into a series of $n(2^{d}-1)$ binary indicator variables (i.e., indicator basis functions), $W$, such that as $n\rightarrow\infty$ the logit of the propensity score function, $logit(e(X))$, can be approximated arbitrarily well by a linear combination of the indicator basis functions written as:
$logit(g(W))=\gamma_{0}+W^{\top}\gamma$ (2)
@@
Throughout this paper, we focus on the application of LASSO models to the baseline covariates, $X$, under the assumption of linearity. However, the methods considered here are generally applicable to nonlinear settings by fitting a LASSO regression on a linear combination of the expanded indicator basis functions, $W$. This has been termed the Highly Adaptive Lasso—a nonparametric machine-learning prediction algorithm that has been shown to have theoretical guarantees on fast convergence rates under mild assumptions [14, 15]. While our focus is not on the Highly Adaptive Lasso, we do consider the application of the Highly Adaptive Lasso in some numerical studies in Section 3. For more on the Highly Adaptive Lasso, see Benkeser & van der Laan [14]. A general overview of the Highly Adaptive Lasso is provided in the Supplemental Appendix.
@@ 2.2 Undersmoothing LASSO PS Models
When using LASSO regression to estimate $e(X)$, the parameter vector, $\beta$, in Equation (1) is estimated based on the following penalized likelihood [4]:
${\cal L}(\beta)=\sum_{i=1}^{n}-A_{i}log(p(A_{i}|X_{i};\beta))-(1-A_{i})log(1-p(A_{i}|X_{i};\beta))+\lambda\sum_{j=1}^{d}\mid\beta_{j}\mid$ (3)
@@
where $\lambda$ is the regularization tuning parameter, $n$ is the sample size, $d$ is the number of parameters in $\beta$, and $p(A_{i}|X_{i};\beta)$ follows the logistic model defined in Equation 1.
@@
If the regularization tuning parameter, $\lambda$, in Equation 3 is chosen appropriately, then as $n\rightarrow\infty$, $\lambda$ will tend to zero and the coefficients within the LASSO model will converge to those in Equation 1. In finite samples, however, LASSO is just an approximation to $e(X)$ and different choices for $\lambda$ provide different LASSO estimators. The optimal choice for $\lambda$ depends on the purpose for which the LASSO model is used. If the goal is to optimize out-of-sample prediction, then $\lambda$ is often chosen using cross-validation, which we will refer to as $\lambda_{CV}$. Ertefaie et al [15] showed, however, that when using LASSO models for PS weighting, less regularization is needed to minimize bias in PS weighted estimators. This is referred to as undersmoothing the LASSO model and corresponds to choosing a value for $\lambda$ that is less than $\lambda_{CV}$ so that the model selects more covariates from $X$ compared to the model using $\lambda_{CV}$.
Undersmoothing can capture more confounder information by including more covariates from $X$ in the model. However, undersmoothing can also harm the overall accuracy of the fitted model by modeling spurious associations in the data (overfitting). Previous work has shown that cross-fitting or using the out-of-fold predictions from the LASSO model can improve properties of effect estimates by reducing problems caused by overfitting [12, 20]. Still, there is a tradeoff with undersmoothing LASSO models for PS estimation. Proper undersmoothing improves efficiency of causal estimators by controlling for more confounder information, but too much undersmoothing will eventually harm the accuracy of the estimated coefficients to a degree where the benefit of including more information in $X$ is outweighed by the cost of a poorly fit model producing unstable predictions [21].
Criteria have been proposed to help investigators properly undersmooth LASSO PS models when using estimators that target populations that are well defined with efficient influence functions that are easy to derive (e.g., inverse probability weighting to estimate the average treatment effect) [15, 22]. However, it remains unclear how to properly undersmooth LASSO PS models when the efficient influence function for the target parameter is difficult to derive or when the target population is not well defined. Here, we consider using balance metrics as a simple and general approach to undersmooth LASSO models for PS weighting.
@@ 2.3 Using Balance Metrics to Undersmooth LASSO PS Models
Metrics for evaluating covariate balance across exposure groups have become standard when evaluating PS models for confounding control [23, 24, 25]. Balance diagnostics have the benefit of being easy to implement and generally applicable. Still, it can be challenging to measure balance on the joint covariate distribution in high-dimensional settings [26]. If $X$ is binary and linearly related to exposure, however, assessing balance on the joint correlation structure is straightforward since balance on the marginal prevalence of the covariates implies balance on the full joint covariate distribution. As discussed previously, large-scale PS analyses in healthcare database studies usually consist of adjusting for a high-dimensional set of sparse binary covariates where linearity assumptions are often reasonable.
@@
  • Largest standardized absolute mean difference. Selects the lambda tuning parameter that minimizes the following after PS weighting: $max[|s_{1}|,|s_{2}|,...,|s_{d}|]$.
  • Average standardized absolute mean difference. Selects the lambda tuning parameter that minimizes the following after PS weighting: $\frac{1}{d}\sum_{k=1}^{d}|s_{k}|$.
When $X$ is not binary, or it is not reasonable to assume linearity in the association between $X$ and the log-odds of exposure, the above metrics are still generally applicable by replacing $X$ with an expanded set of indicator basis functions, $W$, as discussed previously. Since the log-odds of exposure is linear in $W$, it is conjectured that balance on the marginal prevalence of the generated binary indicators implies balance on the joint distribution of the original covariates. It is important to note that for continuous covariates, it is common practice to only balance covariate means, which is restrictive. Balancing the expanded set of indicator basis functions for non-binary covariates allows for stronger, more nonparametric balance guarantees.
While many other balance metrics could be considered, our goal is not to perform an exhaustive evaluation of approaches that could be used to undersmooth LASSO PS weighted estimators. Here, we consider two commonly used balance metrics for evaluating large-scale PS models but emphasize that no single metric is universally best. The best metric will depend on properties of the given study as well as the chosen weighting approach. Consequently, it can be difficult to know which metric best correlates with bias for the study at hand and if the level of balance achieved is adequate to remove bias caused by measured confounders [26]. Therefore, to address this challenge, we suggest supplementing balance diagnostics by using synthetically generated negative control exposures to help detect biased analyses.

### v3

@@ Abstract
The propensity score (PS) is often used to control for large numbers of covariates in high-dimensional healthcare database studies. The least absolute shrinkage and selection operator (LASSO) has become the most widely used tool for fitting large-scale PS models in these settings. LASSO uses $L_{1}$ regularized regression to prevent overfitting by shrinking coefficients toward zero (setting some exactly to zero). The degree of regularization is typically selected using cross-validation to minimize out-of-sample prediction error. Both theory and simulations have shown, however, that when using LASSO models for PS weighting, less regularization is needed to minimize bias in PS weighted estimators. This is referred to as undersmoothing the LASSO model, where the optimal degree of undersmoothing can be derived from the target causal parameter’s efficient influence function. In many settings, however, the efficient influence function is unknown or difficult to derive. Here, we consider the use of balance metrics as a simple and generally applicable approach to select the degree of undersmoothing when the efficient influence function is unknown. Because LASSO models that are tuned using balance metrics alone are not assured to minimize bias in PS weighted estimators—as such metrics are blind to the efficient influence function—we propose a framework to generate synthetic negative control exposures for bias detection. We show that synthetic negative control exposures can identify analyses that likely violate partial exchangeability due to lack of control for measured confounding. Finally, we use a series of numerical studies to investigate the finite sample performance of using balance criteria to undersmooth LASSO PS-weighted estimators, and the use of synthetic negative control exposures to detect biased analyses.
@@ 1 Introduction
Routinely collected healthcare data, including administrative claims and electronic health records, are increasingly being used to generate real‐world evidence on the effects of medical products. However, confounding stemming from nonrandomized exposures remains a fundamental obstacle to effectively utilizing these data sources for real-world evidence generation. To improve confounding control in these settings, several studies have shown that data‐adaptive algorithms can be used to leverage the large volume of information in these data sources to identify and control for large numbers of covariates when estimating nuisance functions for causal inference (i.e., functions that are used to construct estimators for causal effects such as the propensity score (PS) or the conditional outcome mean) [1, 2, 3, 4, 5]. In healthcare database studies, however, outcome events are often rare, making it difficult to empirically identify and model outcome associations for large numbers of covariates. Consequently, the PS, defined as the conditional probability of exposure given a set of covariates, is widely used for large-scale covariate adjustment in these settings [6, 7].
The least absolute shrinkage and selection operator (LASSO) has become the most widely used tool for fitting large-scale PS models [4, 8]. LASSO reduces overfitting by using $L_{1}$ regularized regression to shrink model coefficients toward zero (setting some exactly to zero) [8]. The degree of regularization depends on a penalty term which is typically selected using cross-validation to minimize out-of-sample prediction error. Recent work has shown, however, that when using LASSO models for PS weighting, less regularization is needed to minimize bias in PS weighted estimators [9]. This is referred to as undersmoothing the LASSO model and corresponds to selecting a degree of regularization that is less than the degree of regularization that is selected using cross-validation (discussed further in Section 2).
Ertefaie et al [9] discuss the importance of undersmoothing in the semiparametric estimation setting when using the Highly Adaptive LASSO [10] to construct PS weighted estimators. Building on the work of Benkeser & van der Laan [10], Ertefaie et al [9] showed that if the true PS function can be expressed as a linear combination of an expanded set of indicator basis functions of the covariates (discussed further in Section 2), then a properly undersmoothed LASSO regression that is fitted on the expanded set of basis functions can enable the construction of nonparametric PS weighted estimators that are asymptotically linear with variance converging to the semiparametric efficiency bound. Recent work further suggests that even under the strong structural condition of linearity (where no basis expansion of the covariates is necessary), proper undersmoothing of a LASSO model that is fitted on a linear combination of the covariates can still reduce bias in PS weighted estimators relative to estimators that are tuned using cross-validation [11, 12].
@@
Ertefaie et al [9] showed that the optimal degree of undersmoothing to minimize bias in LASSO PS weighted estimators can be derived from the form of the target causal parameter’s efficient influence function, if that function is available. In practice, however, the efficient influence function is often not known or difficult to derive. Even when the analytic form of the efficient influence function is known, using the efficient influence function to undersmooth the LASSO model requires modeling the conditional outcome mean [9], which can be difficult in studies with rare outcome events. It remains unclear how to best tune LASSO PS weighted estimators when the efficient influence function is difficult to derive or unknown, or when modeling the conditional outcome mean is challenging.
@@
In this work, we discuss challenges in determining how to undersmooth LASSO models for PS weighting. We consider the use of balance metrics as a practical and generally applicable approach to select the degree of undersmoothing when the efficient influence function is unknown or difficult to derive. Because LASSO models that are tuned using balance metrics alone are not assured to minimize bias in PS weighted estimators—as such metrics are blind to the efficient influence function—we propose a framework to generate synthetic negative control exposures for bias detection. We show that if a given LASSO PS weighted analysis does not result in conditional independence between the synthetic exposures and observed outcome within unexposed individuals, then the same analysis is unlikely to satisfy partial exchangeability when applied to the original study population. Finally, we use a series of numerical studies to evaluate the finite sample performance of using balance criteria to undersmooth LASSO PS-weighted estimators, and the use of synthetic negative control exposures to detect biased analyses.
@@ 2.1 Framework & Setup
We assume that the exposure is conditionally exchangeable given $X$, written as
@@
where $\perp\!\!\!\perp$ denotes the conditional independence of random variables. Conditional exchangeability implies no unmeasured confounding. Conditional exchangeability given $X$ implies conditional exchangeability given $e(X)$ [6]. Conditional exchangeability also implies partial exchangeability, written as
@@
We further assume positivity and consistency [16]. Positivity, formally written as $0<e(X)<1$, implies that the true PS function, $e(X)$, is bounded away from 0 and 1. Consistency implies that the observed outcome for each individual is equal to the potential outcome under their observed exposure status, written as $Y^{a}=Y$ for $A=a$.
In addition to the standard assumptions for causal inference described previously, we assume that the logit of the propensity score function, $logit(e(X))$, is linear in $X$, written as:
$logit(e(X))=\beta_{0}+X^{\top}\beta,$ (1)
where $\beta$ is a $d$-dimensional vector of parameters and $\beta_{0}$ is a scalar. Healthcare database studies often involve adjusting for a high-dimensional set of sparse binary covariates where linearity assumptions in the PS model are often reasonable [1, 2, 3, 4, 5]. When the assumption of linearity in the relationship between $X$ and the logit of the propensity score is not reasonable, Benkeser & van der Laan [10] showed that under mild global smoothness assumptions, the baseline covariates, $X$, can be expanded into a series of $n(2^{d}-1)$ binary indicator variables (i.e., indicator basis functions), $W$, such that as $n\rightarrow\infty$ the logit of the propensity score function, $logit(e(X))$, can be approximated arbitrarily well by a linear combination of the indicator basis functions written as:
$logit(g(W))=\gamma_{0}+W^{\top}\gamma,$ (2)
@@
Throughout this paper, we focus on the application of LASSO models to the baseline covariates, $X$, under the strong structural assumption of linearity. However, the methods considered here are generally applicable to nonlinear settings by fitting a LASSO regression on a linear combination of the expanded indicator basis functions, $W$. This has been termed the Highly Adaptive LASSO—a nonparametric machine-learning prediction algorithm that has been shown to have theoretical guarantees on fast convergence rates under mild assumptions [9, 10]. While our focus is not on the Highly Adaptive LASSO, we do consider the application of the Highly Adaptive LASSO in some numerical studies in Section 3. For more on the Highly Adaptive LASSO, see Benkeser & van der Laan [10] and Butzin-Dozier et al [17]. A general overview of the Highly Adaptive LASSO is provided in the Supplemental Appendix.
@@ 2.2 Undersmoothing LASSO PS Models
When using LASSO regression to estimate $e(X)$, the parameter vector $\beta$ in Equation (1) is estimated by minimizing the following penalized negative log-likelihood:
${\cal L}(\beta)=-\sum_{i=1}^{n}\Big[A_{i}\log P(A_{i}\mid X_{i};\beta)+(1-A_{i})\log\big(1-P(A_{i}\mid X_{i};\beta)\big)\Big]+\lambda\sum_{j=1}^{d}|\beta_{j}|.$ (3)
@@
Here, $\lambda$ is the regularization tuning parameter, $n$ is the sample size, $d$ is the number of parameters in $\beta$, and $P(A_{i}|X_{i};\beta)$ follows the logistic model defined in Equation 1.
@@
If the regularization tuning parameter $\lambda=\lambda_{n}$ is chosen such that $\lambda_{n}/n\rightarrow 0$ as $n\rightarrow\infty$, then the LASSO estimator is consistent, and the estimated coefficients within the LASSO model will converge to those in Equation 1 [18, 19]. In finite samples, however, the LASSO solution is just an approximation to $e(X)$ and different choices for $\lambda$ provide different LASSO estimators. The optimal choice for $\lambda$ depends on the purpose for which the LASSO model is used. If the goal is to optimize out-of-sample prediction, then $\lambda$ is often chosen using cross-validation, which we will refer to as $\lambda_{CV}$. Ertefaie et al [9] showed, however, that when using LASSO models for PS weighting, less regularization is needed to minimize bias in PS weighted estimators. This is referred to as undersmoothing the LASSO model and involves choosing a $\lambda$ value that is less than $\lambda_{CV}$, resulting in less shrinkage of the coefficient estimates.
Butzin-Dozier et al [17] explain that undersmoothing is needed for two reasons. First, the $L_{1}$-norm penalization incorporated into the LASSO fitting procedure has the capacity to yield sparse PS models by shrinking some coefficients to exactly zero, which can result in some confounders being excluded from the final model. Second, even for covariates that are selected into the model, LASSO’s $L_{1}$ penalization introduces bias in the estimated coefficients by shrinking them toward zero. While this is benefiical for improving out-of-sample prediction by creating a more stable model, it can be suboptimal for confounding control as it results in the fitted PS model only partially solving the efficient influence functions estimating equations [9, 17]. Undersmoothing addresses these issues by debiasing the PS weighted estimator. Specifically, undersmoothing can enable the construction of asymptotically linear and efficient PS weighted estimators by solving the efficient influence function’s estimating equations and eliminating an asymptotic bias term [9, 17]. It is important to note that while undersmoothing has generally been discussed and applied in the semiparametric estimation setting described by Ertefaie et al [9], recent work suggests that even under the strong structural assumption of linearity (where no basis expansion of the covariates is necessary), undersmoothing the LASSO model can still reduce bias in PS weighted estimators [11, 12].
While undersmoothing can improve the properties of LASSO PS-weighted estimators, it also increases the risk of bias caused by excessive overfitting. Previous work has shown that cross-fitting or using out-of-fold predictions can improve properties of effect estimates by reducing bias caused by overfitting [9, 20, 21, 22, 23]. It is important to note, however, that cross-fitting is not universally required to avoid overfitting biases; its necessity depends on the specific model conditions. For instance, under sufficient sparsity and dimensionality conditions, Chernozhukov et al [24] showed that LASSO-based estimators do not rely on cross-fitting for debiased estimation. Analogously, when the sectional variation norm is bounded—a condition assumed in Ertefaie et al [9]—estimators based on the Highly Adaptive LASSO also do not require cross-fitting for debiased estimation [9]. Nevertheless, Ertefaie et al [9] demonstrated that even under such conditions, cross-fitting can still be beneficial by providing finite-sample improvements. Still, there is a tradeoff with undersmoothing LASSO models for PS estimation. Proper undersmoothing can reduce bias of PS-weighted estimators, but excessive undersmoothing eventually harms the accuracy of the estimated coefficients to a degree where the benefit of undersmoothing is outweighed by the cost of a poorly fit model producing unstable predictions [12].
Criteria have been proposed to help investigators properly undersmooth LASSO PS models when the efficient influence function for the target causal parameter is easy to derive [9, 25]. For settings where the efficient influence function is difficult to derive or unknown, Ertefaie et al [9] proposed a “score-based” criterion to select the degree of undersmoothing in an outcome-blind manner (see section 4.2 of Ertefaie et al for details). While this approach has been shown to perform well across several simulations [9], it is specific to inverse probability weighted estimators. This is because it requires overfitting of the PS model to correspond with an increase in the variance of the weight function to determine when undersmoothing becomes too extreme. Since overfitting need not produce extreme weights for all PS weighted estimators (e.g., matching weights and overlap weights), this criterion is not generally applicable. Here, we consider using balance metrics as a simple and generally applicable approach to undersmooth LASSO models for PS weighting.
@@ 2.3 Using Balance Metrics to Undersmooth LASSO PS Models
Metrics for evaluating covariate balance across exposure groups have become standard when evaluating PS models for confounding control [26, 27, 28]. Balance diagnostics have the benefit of being easy to implement and generally applicable. Still, it can be challenging to measure balance on the joint covariate distribution in high-dimensional settings [29]. If $X$ is linearly related to exposure, however, assessing balance on the joint correlation structure simplifies to assessing balance on the first moment of each covariate (covariate means). As discussed previously, large-scale PS analyses in healthcare database studies usually consist of adjusting for a high-dimensional set of sparse binary covariates where linearity assumptions are often reasonable.
@@
  • Largest standardized absolute mean difference. Selects the $\lambda$ tuning parameter that minimizes the following after PS weighting: $max[|s_{1}|,|s_{2}|,...,|s_{d}|]$.
  • Average standardized absolute mean difference. Selects the $\lambda$ tuning parameter that minimizes the following after PS weighting: $\frac{1}{d}\sum_{k=1}^{d}|s_{k}|$.
When $X$ is not binary, or it is not reasonable to assume linearity in the association between $X$ and the log-odds of exposure, the above metrics are still generally applicable by replacing $X$ with an expanded set of indicator basis functions, $W$, as discussed previously. Since the log-odds of exposure is linear in $W$, it is conjectured that balance on the marginal prevalence of the generated binary indicators implies balance on the joint distribution of the original covariates. Note that for continuous covariates, it is common practice to only balance covariate means, which is restrictive when covariate associations with exposure are nonlinear. Balancing the expanded set of indicator basis functions for non-binary covariates allows for stronger, more nonparametric balance guarantees.
Although balance metrics are simple to implement and generally applicable for tuning LASSO PS models, it is important to emphasize that balance metrics also have inherent limitations [29]. In particular, LASSO models that are tuned using balance metrics alone are not assured to minimize bias in PS weighted estimators as such metrics are blind to the efficient influence function. While balance critera that take into account covariate-outcome associations or simultaneously consider the efficient influence function could potentially address this limitation [30], our objective here is to focus on balance criteria that are applicable to settings where the efficient influence function and covariate-outcome associations are unknown or difficult to estimate. In such settings, no single balance metric is universally best since they are outcome-blind. Consequently, it can be difficult to know which metric best correlates with bias for the study at hand, or if the level of balance achieved is adequate to remove bias caused by measured confounders [29]. Therefore, to address this challenge, we suggest supplementing balance diagnostics by using synthetically generated negative control exposures to help detect biased analyses.